what does x stand for?
4x-(2x=3)=7
how i did it ---
-8x-12x=7
-20x=77/20
x=7/20
thanks

Answer by Alan3354(69443)

4x-(2x=3)=7 Too many equal signs.
-----------------------------------
how i did it ---
-8x-12x=7
-20x=77/20
x=7/20